A sledge loaded with bricks has a total mass of 18.5 kg and is pulled at constant speed by a rope inclined at 19.5° above the horizontal. The sledge moves a distance of 19.3 m on a horizontal surface. The coefficient of kinetic friction between the sledge and surface is 0.500.

(a) What is the tension in the rope?
N

(b) How much work is done by the rope on the sledge?
kJ

(c) What is the mechanical energy lost due to friction?
kJ
Answer #1

(a)


noraml force Fn = mg - T*sintheta


friction force fk = uk*Fn = uk*(m*g - T*sintheta)


along horizontal


acceleration ax = 0

Fnet = 0


Tx - fk = 0


T*costheta - uk*(m*g - T*sintheta) = 0


T*cos19.5 - 0.5*(18.5*9.81- T*sin19.5) = 0

tension T = 81.8 N


=======================


(b)

work done by rope = T*costheta*L = 81.8*cos19.5*19.3 = 1488.2 J

Wrope = 1.50 kJ

=======================

(c)


mechanical energy lost E = fk*L


E = uk*(m*g - T*sintheta)*L

E = 0.5*(18.5*9.81- 81.8*sin19.5)*19.3


E = 1.5 kJ